PART A) A small cork with an excess charge of +5.0 µC is placed 0.19 m from another cork, which carries a charge of −3.2 µC. What is the magnitude of the electric force between the corks? The Coulomb constant is 8.98755 × 109 N · m2 /C 2 . Answer in units of N.

PART B) Is this force attractive or repulsive?

PART C) How many excess electrons are on the negative cork? Answer in units of electrons.

PART D)How many electrons has the positive cork lost? Answer in units of electrons.
Answer #1

A. F = kq1q2/r^2

= (9*10^9)(5*10^-6)(3.2*10^-6) / (0.19)^2

F= 3.99 N Answer

B. Since charges has opposite sings, Attractive Force will be Presenting between them.

Part C.

n=Q2/e

(3.2*10^-6) / (1.6022*10^−19)

= 1.99*10^13 electrons gained. Answer

Part D.

n=Q1/e

(5*10^-6) / (1.6022*10^−19)

= 3.121*10^13 electrons lost Answer

I hope help you !!